OpenStudy (anonymous):

What is the solution of the equation? b+6/5 = 10

OpenStudy (anonymous):

Is only 6 over five, or are both b and 6 over 5

OpenStudy (hba):

b=10-6/5

OpenStudy (hba):

take lcm

OpenStudy (anonymous):

b and 6 are both 5

OpenStudy (anonymous):

b and 6 are both over 5

OpenStudy (anonymous):

(b+6)/5=10 multiply both sides by 5 to get rid of the denominator b+6=50 b=44

OpenStudy (anonymous):